Assume that stock market returns do not resemble a single-index structure. (a) Assume that stock market returns do not...

  1. Assume that stock market returns do not resemble a single-index structure. (a) Assume that stock market returns do not resemble a single-index structure. An investment fund analyzes 2,000 stocks in order to construct a mean-variance efficient portfolio constrained by 2,000 investments. They will need to calculate (N) ____ expected returns, (N)___________ variances of returns and ________ covariances.

  1. Assume that stock market returns do follow a single-index structure. An investment fund analyzes 2,000 stocks in order to construct a mean-variance efficient portfolio constrained by 2,000 investments. They will need to calculate (N)________ estimates of expected returns and (N)________ estimates of sensitivity coefficients to the market factor (betas).
Answer #1

Ans 1):Assume that stock market returns do not resemble a single-index structure. (a) Assume that stock market returns do not resemble a single-index structure. An investment fund analyzes 2,000 stocks in order to construct a mean-variance efficient portfolio constrained by 2,000 investments. They will need to calculate (N) 2000 expected returns, (N)2000 variances of returns and 1,990,000.covariances.

Explanation:

  1. The expected returns of each of the 2000 securities must be calculated so as to get the detail variation of each stock i.e. to include all the 2000 stock.
  2. The 2000 variances around these stock returns must be calculated.
  3. We will calculate the covariance by the below formula

Covariance=(n2 − n)/2

Here N=2000

Covariance=(4,000,000 − 150)/2 = 1,990,000.

Ans 2 ): Assume that stock market returns do follow a single-index structure. An investment fund analyzes 2,000 stocks in order to construct a mean-variance efficient portfolio constrained by 2,000 investments. They will need to calculate (N)2000 estimates of expected returns and (N)2000 estimates of sensitivity coefficients to the market factor (betas).

Explanation:

For a single-index model, N=2000, expected returns and N=2000 sensitivity coefficients to the market factor (betas) must be estimated.